# Hot, Photoionized X-ray Gas in Two Luminous Type 2 Quasars: Chandra-HST Evidence for a Wind-Driven Sequence

### v1

@@ 4.2 Broad- and Narrow-Band Imaging
Figures 3 and 4 show the resulting PSF-subtracted images, smoothed with a Gaussian kernel () and displayed with asinh scaling. White contours trace the [O III] emission (fischer_hubble_2018), allowing for direct comparison between the X-ray and optical morphologies.
@@
The full-band (0.3-8 keV) emission extends to (4.1 kpc). The soft X-rays dominate, showing clumpy structures in multiple directions and two distinct peaks are evident: one located 0.2′′ (410 pc) east of the centroid, likely associated with the peaks seen in the Neon and Magnesium band images, and another 0.2′′ (410 pc) southwest of the centroid, consistent with the peak observed in the Silicon band. The soft-band morphology closely follows the [O III] contours, with the brightest X-ray knots coincident with peaks seen in [O III]. The hard-band emission is less extended (, 2.9 kpc) but shows similar correspondence with the inner [O III] structures. In the narrow bands, the Neon band has a low signal-to-noise ratio (15 counts within 3.5′′) but still shows extended emission peaking 0.35′′ (715 pc) northeast of the nucleus. The Magnesium band (28 counts) shows a clumpy feature 0.6′′ (1.2 kpc) southwest of the nucleus, while the Silicon band (25 counts) peaks west of the nucleus, coincident with the brightest [O III] peak.
@@
5 X-ray Gas Modeling
5.1 Photoionization Modeling
We construct photoionization grids using Cloudy version C.23 (chatzikos_2023_2023). The predicted line spectrum depends on the shape of the ionizing SED, the gas density (), column density (), distance from the ionizing source (), and the adopted chemical abundances. The models are parameterized by the dimensionless ionization parameter
@@ 5.1 Photoionization Modeling
Following Section 3.2, we adopt an SED of the form , with below 13.6 eV, between 13.6 eV and 0.5 keV, from 0.5–10 keV, and from 10–100 keV, with a high-energy cutoff at 100 keV. The ionizing luminosities are taken from trindade_falcao_hubble_2021-1, based on the NuSTAR analysis of gandhi_nustar_2014: photons s-1 for FIRST J120041 and photons s-1 for 2MASX J130038 (but see Section 5.2). Chemical abundances are set to solar, consistent with our earlier work.
Because the emitting regions lie at distances of tens of parsecs or more, we adopt an open-slab geometry, following previous studies of extended photoionized gas (e.g., kraemer_physical_2015; revalski_quantifying_2018; kraemer_mass_2020; trindade_falcao_hubble_2021-1).
5.2 Modeled Line Luminosities
@@
At a fixed distance from the nucleus, , the only free parameters in Equation 4 are and . For each model in the Cloudy grid, we therefore select a range of values that reproduce the observed emission lines and solve for the physically consistent density. We adopt , the characteristic average radius of the extraction region, for both quasars.
@@
To identify the best-fitting model at each location, we compare the modeled and observed line ratios for all fractional combinations of LOW and HIGH components; an ideal match yields a ratio of unity. The ionizing luminosities are uncertain by a factor of 4 (trindade_falcao_hubble_2021-1). For FIRST J120041, the published value reproduces the observed luminosities, whereas for 2MASX J130038 a value larger by a factor of 3.8 is required, still within the allowed range.
Physically, the models are constrained by two requirements: (1) the emitting slab thickness, , cannot exceed the available gas column at that radius, and (2) the emitting area cannot exceed the maximum geometric surface subtended by the source (see also Section 3.2).
5.3 Mass of Ionized Gas
We estimate the total mass of the photoionized X-ray–emitting gas by combining the observed line luminosities with Cloudy photoionization models (Section 5.1). Following kraemer_physical_2015; kraemer_mass_2020, the mass of gas at radius is
@@ 5.3 Mass of Ionized Gas
Observed fluxes are measured using srcflux121212https://cxc.cfa.harvard.edu/ciao/ahelp/srcflux.html and converted to luminosities using distances from fischer_hubble_2018. Table 4 lists the observed and modeled line luminosities within for each target.
@@
where and are the modeled fluxes of each component. We solve for the emitting areas by fitting the Neon, Magnesium, and Silicon bands simultaneously, using the SciPy differential_evolution131313https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.differential_evolution.html algorithm with geometric constraints as bounds. Uncertainties are estimated by perturbing the observed luminosities within their measurement errors and repeatedly re-optimizing in a Monte Carlo framework.
5.4 Comparison with Optical Ionized [O III] Gas
We now compare the mass of the X-ray–emitting gas to that of the cooler ionized phase traced by [O III]. For simplicity, we assume that each phase is uniformly distributed within its respective extraction region. Under this assumption, a direct comparison requires accounting for the fact that the two phases occupy different volumes. We therefore define a volume-normalized X-ray mass
where and are the emitting volumes of the [O III] and X-ray phases. The ratio thus quantifies the relative contributions of the hot and warm ionized gas under a common geometric framework.
In reality, local gas density is expected to decrease with radius (e.g., revalski_quantifying_2022). Moreover, the [O III] phase, having lower ionization and higher density than the X-ray–emitting gas, requires a correspondingly smaller volume to reproduce its observed luminosity. These effects imply that assuming uniform density and unity filling factor tends to overestimate the emitting volumes, especially for the X-ray gas. Consequently, the volume-normalized mass ratios derived below should be regarded as conservative upper limits, particularly in systems where the X-ray–emitting gas extends over a substantially larger region than the [O III] gas.
@@ 5.4 Comparison with Optical Ionized [O III] Gas
Within (2 kpc), the low- and high-ionization X-ray components contain and , respectively. The total X-ray mass is therefore dominated by the low-ionization phase: . The [O III] outflow contains (trindade_falcao_hubble_2021-1) within (1 kpc). After accounting for the different emitting volumes (Equation 7), the resulting volume-normalized mass ratio is .
@@
Within (1.6 kpc), the X-ray–emitting gas contains . The [O III] outflow contains within (0.16 kpc). After correcting for the different emitting volumes, the resulting volume-normalized mass contrast is .
For comparison, nearby Seyfert galaxies host significantly smaller X-ray gas reservoirs. NGC 1068 contains within the inner pc, combining low- and high-ionization components (kraemer_physical_2015), while NGC 4151 has , within pc (kraemer_mass_2020). In, the X-ray gas masses of FIRST J120041 and 2MASX J130038 exceed these values by factors of 300-800 (2.5-3 dex), though measured within volumes 30-8000 times larger.
@@
6 The Role of X-ray Gas
6.1 [O III] Gas Kinematics
To investigate whether the X-ray and optical phases are dynamically correlated, we compare the X-ray surface brightness with the kinematics of the ionized [O III] gas (fischer_hubble_2018). In that study, the authors defined three kinematic classes based on the [O III] velocity centroids relative to systemic and line widths, which we adopt here:
@@ 6 The Role of X-ray Gas
Non-disturbed rotation: [O III] lines with centroid velocities km s-1 and FWHM km s-1, consistent with ordered disk rotation (see also bellocchi_vltvimos_2013; ramos_almeida_infrared_2017).
We use the dmextract141414https://cxc.cfa.harvard.edu/ciao/ahelp/dmextract.html task to measure the surface brightness of the 0.3-3 keV soft X-ray emission along rectangular extraction strips aligned with the HST/STIS slit positions from fischer_hubble_2018. Each strip spans in length and is divided into nine bins, matching the angular resolution of Chandra. Figures 6 and 7 show, respectively, the extraction geometry overlaid on the 0.3-3 keV Chandra images (top left) and the STIS slit over the HST [O III] images (top right). The bottom panels compare the spatially resolved X-ray surface brightness profiles (orange) with the [O III] velocity centroids (purple, bottom left) and [O III] FWHM values (green, bottom right) for both quasars.
@@
The soft X-ray surface brightness profile confirms that the emission peaks off-nucleus, with a clear enhancement to the east of the nucleus ( pc). Out to at least kpc in that direction, the X-ray surface brightness remains higher than at the nuclear position. Comparing this distribution to the [O III] outflowing kinematics (purple points above the reference line in the bottom-left panel), we find that regions of enhanced X-ray surface brightness (40 counts arcsec-2) coincide with locations where the [O III] gas shows the highest velocities (-750 km s-1) and the broadest line widths (FWHM up to 1700 km s-1). This spatial correspondence shows that the brightest, hot X-ray-emitting gas is co-spatial with the outflowing [O III] component, suggesting a direct interaction between the two phases. While the [O III] flux is higher in the right side (western) emission knot, the outflow velocities and FWHMs are greater on the left (eastern) side, exactly where the soft X-ray peak happens. This further supports the interpretation that the enhanced X-ray emission pinpoints regions of strong coupling between the hot ionized and the cooler [O III] outflow detected by fischer_hubble_2018.
@@
The soft X-ray surface brightness profile also peaks off-nucleus, showing enhancements on both sides of the nucleus. Out to 1 kpc in either direction along the slit, the extended emission has higher surface brightness than the nuclear region. Although the per-bin X-ray surface brightness (orange points in Figure 7) is higher than in FIRST J120041, the [O III] kinematics in this quasar (purple points, bottom-left panel) are consistent with ordered rotation, i.e., not-outflowing (fischer_hubble_2018). The centroid velocities remain below the threshold for outflowing gas (purple line), and the FWHM values are generally below the disturbed gas threshold (green line), except for a localized broadened component pc northwest of the nucleus. This broadened region coincides with the base of the linear [O III] structure observed near the nucleus. In the X-ray image, an enhancement appears at the end of this structure (see also Figure 4) and another to the southeast, at the location of the prominent [O III] arc. The combination of high FWHM (250 km s-1) but low centroid velocity (300 km s-1) at these position falls into the “disturbed gas” category of fischer_hubble_2018, suggesting that small-scale turbulence or nascent outflow activity could be arising from the linear [O III] feature. Despite its relatively high X-ray surface brightness, no [O III] outflows are detected in this system, and Figure 7 shows no clear spatial correlation between the X-ray enhancements and the [O III] velocity or line-width distributions.
6.2 A Relativistic Hot Wind?
@@
The spectrum of 2MASX J130038 shows a strong neutral Fe K line at 6.4 keV (EW = 228 eV) together with an emission feature at keV (Figures 1, 2). While the mytl component (shown in red) successfully reproduces both the neutral Fe K line at keV and Fe K at keV, an additional Gaussian component centered at keV is required to adequately model the data (significant at 1.4, or 93% confidence).
If interpreted as blueshifted Fe XXVI Ly line (rest energy 6.97 keV), this feature would trace highly-ionized gas either in the immediate vicinity of the nucleus or within a fast nuclear outflow. The observed energy shift corresponds to a velocity of approximately 7,000-8,000 km s-1, consistent with photoionized winds or disk-driven outflows frequently detected in luminous AGNs (e.g., tombesi_evidence_2010; tombesi_wind_2015).
@@
An alternative identification such as Fe K fluorescence from neutral iron (rest energy 7.06 keV) would require a much larger flux, well above the expected 10-15% of Fe K flux (molendi_iron_2003; bianchi_soft_2006). Moreover, the mytl component already reproduces the expected Fe K line, whose flux is 16% of the fitted Fe K line intensity, as shown in Figures 1 and 2. Emission from Ni K (rest energy 7.47 keV) lies too far from the measured centroid to provide a satisfactory explanation, though this possibility cannot be completely ruled out given the current data.
@@
We therefore interpret the 7.14 keV feature as tentative evidence for Fe XXVI Ly emission arising from a hot, blueshifted nuclear wind. In contrast to the absorption-dominated UFOs commonly observed in quasars (e.g., tombesi_discovery_2010), this candidate outflow is detected in emission, implying a geometry and covering factor that allow significant reprocessing of the central radiation field. Deeper hard X-ray observations will be required to confirm its presence and to better constrain its physical conditions and geometry. We discuss the implications of this putative hot wind in the context of a proposed evolutionary sequence for quasars in Section 7.
7 An Evolutionary Path for Quasars

### v2

@@ IV.2 Broad- and Narrow-Band Imaging
Figures 3 and 4 show the resulting PSF-subtracted images, smoothed with a Gaussian kernel () and displayed with asinh scaling. White contours trace the [O III] emission from archival HST narrow-band imaging (Fischer et al., 2018). The contours start at above the background and increase in powers of two (i.e., ), such that the outermost contour marks the lowest significant [O III] emission and successive inner contours correspond to progressively higher emission-line surface brightness. This allows for a direct, quantitative comparison between the X-ray and optical morphologies.
@@
The full-band (0.3-8 keV) emission extends to (4.1 kpc). The soft X-rays dominate, showing clumpy structures in multiple directions and two distinct peaks are evident: one located 0.2′′ (410 pc) east of the centroid, likely associated with the peaks seen in the Neon and Magnesium band images, and another 0.2′′ (410 pc) southwest of the centroid, consistent with the peak observed in the Silicon band. The soft-band morphology is spatially consistent with the [O III] emission, with the brightest X-ray knots preferentially located within the higher-significance [O III] contours. The hard-band emission is less extended (, 2.9 kpc) but shows similar correspondence with the inner [O III] structures. In the narrow bands, the Neon band has a low signal-to-noise ratio (15 counts within 3.5′′) but still shows extended emission peaking 0.35′′ (715 pc) northeast of the nucleus. The Magnesium band (28 counts) shows a clumpy feature 0.6′′ (1.2 kpc) southwest of the nucleus, while the Silicon band (25 counts) peaks west of the nucleus, coincident with the brightest [O III] peak.
@@
In all panels, the color scale represents the X-ray counts per pixel, displayed using a sinh-1 stretch to enhance low-surface-brightness emission.
V X-ray Gas Modeling
V.1 Photoionization Modeling
We construct photoionization grids using Cloudy version C.23 (Chatzikos et al., 2023). The predicted line spectrum depends on the shape of the ionizing SED, the gas density (), column density (), distance from the ionizing source (), and the adopted chemical abundances. The models are parameterized by the dimensionless ionization parameter
@@ V.1 Photoionization Modeling
Following Section III.2, we adopt an SED of the form , with below 13.6 eV, between 13.6 eV and 0.5 keV, from 0.5–10 keV, and from 10–100 keV, with a high-energy cutoff at 100 keV. The ionizing luminosities are taken from Trindade Falcão et al. (2021b), based on the NuSTAR analysis of Gandhi et al. (2014): photons s-1 for FIRST J120041 and photons s-1 for 2MASX J130038 (but see Section V.2). Chemical abundances are set to solar, motivated by photoionization studies of AGN NLRs, which generally indicate modestly supersolar metallicities. In particular, Groves et al. (2004) inferred nitrogen abundances of approximately twice solar in NLR gas; since nitrogen is a secondary element whose abundance scales as (Talbot and Arnett, 1973), this corresponds to an overall metallicity of approximately 1.4, which we adopt here (e.g., Trindade Falcão et al., 2021b).
V.2 Modeled Line Luminosities
@@
At a fixed distance from the nucleus, , the only free parameters in Equation 4 are and . For each model in the Cloudy grid, we therefore select a range of values that reproduce the observed emission lines and solve for the physically consistent density.
We adopt , the characteristic average radius of the extraction region, for both quasars. Although the X-ray emission is spatially resolved in surface brightness (Section IV; Figure 5), the radial profiles are based on total, band-integrated counts and do not isolate individual emission lines. Determining gas densities and masses requires spectrally decomposed line luminosities and constraints on the ionization parameter in each radial bin. The photon statistics within individual annuli are insufficient to perform such independent spectral modeling as a function of radius. We are therefore restricted to a single, spatially integrated spectral measurement and cannot robustly model multiple radial zones or resolve a density gradient. The chosen radius represents the region enclosing a substantial fraction of the soft X-ray emission and provides a physically reasonable average distance for the line-emitting gas. We note that adopting a larger characteristic radius would decrease the inferred gas density as at fixed ionization parameter (see Equation 4). However, to reproduce the observed line luminosities, the emitting area must increase proportionally as . As a result, the total gas mass, which scales as (see Equation 5), is largely insensitive to the assumed radius. Consequently, even if the line-emitting gas is distributed over larger radii, the derived mass estimates remain robust within the uncertainties.
@@
To identify the best-fitting model at each location, we compare the modeled and observed line ratios for all fractional combinations of LOW and HIGH components; an ideal match yields a ratio of unity. The ionizing luminosities are uncertain by a factor of 4 (Trindade Falcão et al., 2021b). For FIRST J120041, the published value reproduces the observed luminosities, whereas for 2MASX J130038 a value larger by a factor of 3.8 is required, still within the allowed range.
Physically, the models are constrained by two requirements: (1) the emitting slab thickness, , cannot exceed the available gas column at that radius, and (2) the emitting area cannot exceed the maximum geometric surface subtended by the source (see also Section III.2).
V.3 Mass of Ionized Gas
We estimate the total mass of the photoionized X-ray-emitting gas by combining the observed line luminosities with Cloudy photoionization models (Section V.1). Following Kraemer et al. (2015, 2020), the mass of gas at radius is
@@ V.3 Mass of Ionized Gas
Observed fluxes are measured using srcflux111111https://cxc.cfa.harvard.edu/ciao/ahelp/srcflux.html and converted to luminosities using distances from Fischer et al. (2018). Table 4 lists the observed and modeled line luminosities within for each target.
@@
where and are the modeled fluxes of each component.
We solve for the emitting areas by fitting the Neon, Magnesium, and Silicon bands simultaneously, using the SciPy differential_evolution121212https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.differential_evolution.html algorithm with geometric constraints as bounds. The differential_evolution method is a global, population-based optimization algorithm that explores the full parameter space without requiring initial guesses or assuming local convexity of the solution. This approach is well suited for our problem, where the emitting areas are subject to physical bounds and the parameter space may contain multiple local minima. Uncertainties are estimated by perturbing the observed luminosities within their measurement errors and repeatedly re-optimizing in a Monte Carlo framework.
V.4 Comparison with Optical Ionized [O III] Gas
We now compare the mass of the X-ray-emitting gas to that of the cooler ionized phase traced by [O III]. For simplicity, we assume that each phase is uniformly distributed within its respective extraction region. Under this assumption, a direct comparison requires accounting for the fact that the two phases occupy different volumes. We therefore define a volume-normalized X-ray mass
where and are the emitting volumes of the [O III] and X-ray phases, respectively.
The emitting volumes are estimated assuming spherical geometries whose radii correspond to the extraction regions used for each phase. Specifically, is computed assuming a sphere of radius for both targets, matching the X-ray spectral extraction region, while is computed assuming spheres with radii equal to the extent of the [O III] outflowing regions ( for FIRST J120041 and for 2MASX J130038). The volume ratio in Equation 7 therefore reflects the different physical scales probed by the X-ray and [O III] measurements, rather than a pixel-by-pixel comparison within the common field of view, and allows a direct comparison of the relative contributions of the hot and warm ionized gas under a common geometric framework.
In reality, local gas density is expected to decrease with radius (e.g., Revalski et al., 2022). Moreover, the [O III] phase, having lower ionization and higher density than the X-ray-emitting gas, requires a correspondingly smaller volume to reproduce its observed luminosity. These effects imply that assuming uniform density and unity filling factor tends to overestimate the emitting volumes, especially for the X-ray gas. Consequently, the volume-normalized mass ratios derived below should be regarded as conservative upper limits, particularly in systems where the X-ray-emitting gas extends over a substantially larger region than the [O III] gas.
@@ V.4 Comparison with Optical Ionized [O III] Gas
Within (2 kpc), the low- and high-ionization X-ray components contain and , respectively. The total X-ray mass is therefore dominated by the low-ionization phase: . The [O III] outflow contains (Trindade Falcão et al., 2021b) within (1 kpc). After accounting for the different emitting volumes defined by the respective extraction radii (Equation 7), the resulting volume-normalized mass ratio is .
@@
Within (1.6 kpc), the X-ray-emitting gas contains . The [O III] outflow contains within (0.16 kpc). After correcting for the different emitting volumes, the resulting volume-normalized mass contrast is .
For comparison, nearby Seyfert galaxies host significantly smaller X-ray gas reservoirs. NGC 1068 contains within the inner pc, combining low- and high-ionization components (Kraemer et al., 2015), while NGC 4151 has , within pc (Kraemer et al., 2020). The X-ray gas masses of FIRST J120041 and 2MASX J130038 exceed these values by factors of 300-800 (2.5-3 dex), though measured within volumes 30-8000 times larger.
@@
VI The Role of X-ray Gas
VI.1 [O III] Gas Kinematics
To investigate whether the X-ray and optical phases are dynamically correlated, we compare the X-ray surface brightness with the kinematics of the ionized [O III] gas (Fischer et al., 2018). In that study, the authors defined three kinematic classes based on the [O III] velocity centroids relative to systemic and line widths, which we adopt here:
@@ VI The Role of X-ray Gas
We use the dmextract131313https://cxc.cfa.harvard.edu/ciao/ahelp/dmextract.html task to measure the surface brightness of the 0.3-3 keV soft X-ray emission along rectangular extraction strips aligned with the HST/STIS slit positions from Fischer et al. (2018). Each strip spans in length and is divided into nine bins, matching the angular resolution of Chandra. Figures 6 and 7 show, respectively, the extraction geometry overlaid on the 0.3-3 keV Chandra images (top left) and the STIS slit over the HST [O III] images (top right). The bottom panels compare the spatially resolved X-ray surface brightness profiles (orange) with the [O III] velocity centroids (purple, bottom left) and [O III] FWHM values (green, bottom right) for both quasars.
@@
The soft X-ray surface brightness profile confirms that the emission peaks off-nucleus, with a clear enhancement to the east of the nucleus ( pc). Out to at least kpc in that direction, the X-ray surface brightness remains higher than at the nuclear position. Comparing this distribution to the [O III] outflowing kinematics (purple points above the reference line in the bottom-left panel), we find that regions of enhanced X-ray surface brightness (40 counts arcsec-2) coincide with locations where the [O III] gas shows the highest velocities (-750 km s-1) and the broadest line widths (up to 1700 km s-1). This spatial correspondence shows that the brightest, hot X-ray-emitting gas is co-spatial with the outflowing [O III] component, suggesting a direct interaction between the two phases. While the [O III] flux is higher in the right side (western) emission knot, the outflow velocities and FWHMs are greater on the left (eastern) side, exactly where the soft X-ray peak happens. This further supports the interpretation that the enhanced X-ray emission pinpoints regions of strong coupling between the hot ionized and the cooler [O III] outflow detected by Fischer et al. (2018).
@@
The soft X-ray surface brightness profile also peaks off-nucleus, showing enhancements on both sides of the nucleus. Out to 1 kpc in either direction along the slit, the extended emission has higher surface brightness than the nuclear region. Although the per-bin X-ray surface brightness (orange points in Figure 7) is higher than in FIRST J120041, the [O III] kinematics in this quasar (purple points, bottom-left panel) are consistent with ordered rotation, i.e., not-outflowing (Fischer et al., 2018). The centroid velocities remain below the threshold for outflowing gas (purple line), and the FWHM values are generally below the disturbed gas threshold (green line), except for a localized broadened component pc northwest of the nucleus. This broadened region coincides with the base of the linear [O III] structure observed near the nucleus. In the X-ray image, an enhancement appears at the end of this structure (see also Figure 4) and another to the southeast, at the location of the prominent [O III] arc. The combination of high FWHM (250 km s-1) but low centroid velocity (300 km s-1) at these position falls into the “disturbed gas” category of Fischer et al. (2018), suggesting that small-scale turbulence or nascent outflow activity could be arising from the linear [O III] feature. Despite its relatively high X-ray surface brightness, no [O III] outflows are detected in this system, and Figure 7 shows no clear spatial correlation between the X-ray enhancements and the [O III] velocity or line-width distributions.
VI.2 A Relativistic Hot Wind?
@@
The spectrum of 2MASX J130038 shows a strong neutral Fe K line at 6.4 keV (EW = 228 eV) together with an additional emission feature at keV (Figures 1, 2). While the mytl component successfully reproduces both the neutral Fe K line at keV and the associated Fe K line at keV, residual emission remains at higher energies, motivating the inclusion of an additional Gaussian component.
When adding a Gaussian line with width keV and free centroid energy and normalization to the best-fit broadband continuum model derived in Sections III.1 and III.2, with the continuum parameters held fixed, the fit improves by Cstat for two additional free parameters. Under the asymptotic approximation, this corresponds to a -value of (i.e., ). We therefore regard this feature as marginal but potentially meaningful, and treat it with appropriate caution in the discussion below.
If interpreted as blueshifted Fe XXVI Ly emission (rest energy 6.97 keV), the observed centroid implies an outflow velocity of - km s-1. Such velocities are comparable to those inferred for highly ionized nuclear winds in luminous AGNs, which are typically detected in absorption (e.g., Tombesi et al., 2010, 2015). Given the modest statistical significance of the feature, we refrain from drawing strong conclusions and consider this identification suggestive rather than definitive.
@@
Alternative identifications are less satisfactory. Fe K fluorescence from neutral iron (rest energy 7.06 keV) would require a flux well above the expected 10-15% of the Fe K line (Molendi et al., 2003; Bianchi et al., 2006). Moreover, the mytl component already reproduces the expected Fe K emission, with a fitted flux 16% of the Fe K intensity (Figures 1, 2). Emission from Ni K (rest energy 7.47 keV) lies too far from the measured centroid to provide a satisfactory explanation, although this possibility cannot be entirely excluded given the current data quality.
@@
We therefore note the presence of a possible high-energy emission feature consistent with Fe XXVI Ly from a fast, highly ionized nuclear wind, while emphasizing that deeper hard X-ray observations are required to confirm its nature and physical origin. In contrast to the ultra-fast outflows commonly reported in quasars, which are detected in absorption (e.g., Tombesi et al., 2010, 2015), this candidate outflow would be observed in emission, potentially implying a geometry and covering factor that allow substantial reprocessing of the nuclear radiation field. We discuss the implications of such a scenario in the context of quasar evolution in Section VII.
VII An Evolutionary Path for Quasars